Administration has decided to install one at the fifty-yard line of the football stadium hoping to lure people in from the boulevard on game day. Please calculate the pressure in a feed pipe needed to eject a stream of water 30m into the air given atmospheric pressure outside the pipe and standard values for gravity and water density. Also, what is the exit velocity? (I know this looks like a fluids problem...and it is...but it is rooted in thermodynamics ) solve it in thermodynamics way.
Answer #1

exit velocity = sqrt(2 * 9.8 * 30)

                    = 24.248 m/s

=> the pressure in a feed pipe needed = 1/2 * 1000 * 24.2482

                                                                = 2.94 * 105 Pa